# Magpie: Alignment Data Synthesis from Scratch by Prompting Aligned LLMs with Nothing

### v1

@@ Abstract
High-quality instruction data is critical for aligning large language models (LLMs). Although some models, such as Llama-3-Instruct, have open weights, their alignment data remain private, which hinders the democratization of AI. High human labor costs and a limited, predefined scope for prompting prevent existing open-source data creation methods from scaling effectively, potentially limiting the diversity and quality of public alignment datasets. Is it possible to synthesize high-quality instruction data at scale by extracting it directly from an aligned LLM? We present a self-synthesis method for generating large-scale alignment data named Magpie. Our key observation is that aligned LLMs like Llama-3-Instruct can generate a user query when we input only the left-side templates up to the position reserved for user messages, thanks to their auto-regressive nature. We use this method to prompt Llama-3-Instruct and generate 4 million instructions along with their corresponding responses. We perform a comprehensive analysis of the extracted data and select 300K high-quality instances. To compare Magpie data with other public instruction datasets (e.g., ShareGPT, WildChat, Evol-Instruct, UltraChat, OpenHermes, Tulu-V2-Mix), we fine-tune Llama-3-8B-Base with each dataset and evaluate the performance of the fine-tuned models. Our results indicate that in some tasks, models fine-tuned with Magpie perform comparably to the official Llama-3-8B-Instruct, despite the latter being enhanced with 10 million data points through supervised fine-tuning (SFT) and subsequent feedback learning. We also show that using Magpie solely for SFT can surpass the performance of previous public datasets utilized for both SFT and preference optimization, such as direct preference optimization with UltraFeedback. This advantage is evident on alignment benchmarks such as AlpacaEval, ArenaHard, and WildBench, and importantly, it is achieved without compromising performance on reasoning tasks like MMLU-Redux, despite the alignment tax.
@@ 1 Introduction
Figure 1: This figure illustrates the process of self-synthesizing instruction data from aligned LLMs (e.g., Llama-3-8B-Instruct) to create a high-quality instruction dataset. In Step 1, we input only the pre-query template into the aligned LLM and generate an instruction along with its response using auto-regressive generation. In Step 2, we use a combination of a post-query template and another pre-query template to wrap the instruction from Step 1, prompting the LLM to generate the query for the second turn. This completes the construction of the instruction dataset. Magpie efficiently generates diverse and high-quality instruction data. Our experimental results show that Magpie outperforms other public datasets for aligning Llama-3-8B-base.
@@
Large language models (LLMs) such as GPT-4 [1] and Llama-3 [40] have become integral to AI applications due to their exceptional performance on a wide array of tasks by following instructions. The success of LLMs is heavily reliant on the data used for instruction fine-tuning, which equips them to handle a diverse range of tasks, including those not encountered during training. The effectiveness of this instruction tuning depends crucially on access to high-quality instruction datasets. However, the alignment datasets used for fine-tuning models like Llama-3-Instruct are typically private, even when the model weights are open, which impedes the democratization of AI and limits scientific research for understanding and enhancing LLM alignment.
@@
To address the challenges in constructing such datasets, researchers have developed two main approaches. The first type of method involves human effort to generate and curate instruction data [14, 26, 64, 65, 66], which is both time-consuming and labor-intensive [37]. In contrast, the second type of method uses LLMs to produce synthetic instructions [16, 31, 46, 47, 53, 55, 58, 59]. Although these methods reduce human effort, its success heavily depends on prompt engineering and the careful selection of initial seed questions. The diversity of synthetic data tends to decrease as the dataset size grows. Despite ongoing efforts, the scalable creation of high-quality and diverse instruction datasets continues to be a challenging problem.
@@
Is it possible to synthesize high-quality instructions at scale by directly extracting data from advanced aligned LLMs themselves? A typical input to an aligned LLM contains three key components: the pre-query template, the query, and the post-query template. For instance, an input to Llama-2-chat could be “[INST] Hi! [/INST]”, where [INST] is the pre-query template and [/INST] is the post-query template. These templates are predefined by the creators of the aligned LLMs to ensure the correct prompting of the models. We observe that when we only input the pre-query template to aligned LLMs such as Llama-3-Instruct, they self-synthesize a user query due to their auto-regressive nature. Our preliminary experiments indicate that these random user queries are of high quality and great diversity, suggesting that the abilities learned during the alignment process are effectively utilized.
@@
Based on these findings, we developed a self-synthesis method to construct high-quality instruction datasets at scale, named Magpie (as illustrated in Figure 1). Unlike existing methods, our approach does not rely on prompt engineering or seed questions. Instead, it directly constructs instruction data by prompting aligned LLMs with a pre-query template for sampling instructions. We applied this method to the Llama-3-8B-Instruct and Llama-3-70B-Instruct models, creating two instruction datasets: Magpie-Air and Magpie-Pro, respectively.
@@
Our Magpie-Air and Magpie-Pro datasets were created using 206 and 614 GPU hours, respectively, without requiring any human intervention or API access to production LLMs like GPT-4. Additionally, we generated two multi-turn instruction datasets, Magpie-Air-MT and Magpie-Pro-MT, which contain sequences of multi-turn instructions and responses. The statistics and advantages of our instruction datasets compared to existing ones are summarized in Table 1. We perform a comprehensive analysis of the generated data, allowing practitioners to filter and select data instances from these datasets for fine-tuning according to their particular needs.
@@
To compare Magpie data with other public instruction datasets (e.g., ShareGPT [10], WildChat [64], Evol Instruct [58], UltraChat [16], OpenHermes [49], Tulu V2 Mix [24]) and various preference tuning strategies with UltraFeedback [13], we fine-tune the Llama-3-8B-Base model with each dataset and assess the performance of the resultant models on LLM alignment benchmarks such as AlpacaEval 2 [33], Arena-Hard [32], and WildBench [34]. Our results show that models fine-tuned with Magpie achieve superior performance, even surpassing the official Llama-3-8B-Instruct model on AlpacaEval, which was fine-tuned with over 10 million data points for supervised fine-tuning (SFT) and follow-up feedback learning. Not only does Magpie excel in SFT alone compared to prior public datasets that incorporate both SFT and preference optimization (e.g., direct preference optimization with UltraFeedback [13]), but it also delivers the best results when evaluated against six baseline instruction datasets and four preference tuning methods (DPO [44], IPO [2], KTO [19], and ORPO [23] with the UltraFeedback dataset). These findings show the exceptional quality of instruction data generated by Magpie, enabling it to outperform even the official, extensively optimized LLMs.
Table 1: Statistics of instruction datasets generated by Magpie compared to other instruction datasets. Tokens are counted using the tiktoken library [42].
Synthetic Alpaca [47] 52K 1 Low text-davinci-003 67.38±54.88 3.5M
Evol Instruct [58] 143K 1 Low ChatGPT 473.33±330.13 68M
UltraChat [16] 208K 3.16 Low GhatGPT 376.58±177.81 238M
Human Dolly [14] 15K 1 High ChatGPT 94.61±135.84 1.42M
ShareGPT [66] 112K 4.79 High ChatGPT 465.38±368.37 201M
WildChat [64] 652K 2.52 High GPT-3.5 & GPT-4 727.09±818.84 852M
LMSYS-Chat-1M [65] 1M 2.01 High Mix 260.37±346.97 496M
Mixture Deita [38] 9.5K 22.02 - Mix 372.78±182.97 74M
OpenHermes [49] 243K 1 - Mix 297.86±258.45 72M
Tulu V2 Mixture [24] 326K 2.31 - Mix 411.94±447.48 285M
Magpie Llama-3-Magpie-Air 3M 1 No Llama-3-8B 426.39±217.39 1.28B
Llama-3-Magpie-Air-MT 300K 2 No Llama-3-8B 610.80±90.61 366M
Llama-3-Magpie-Pro 1M 1 No Llama-3-70B 478.00±211.09 477M
Llama-3-Magpie-Pro-MT 300K 2 No Llama-3-70B 554.53±133.64  [^3] M
2 Magpie: A Scalable Method to Synthesize Instruction Data
Overview of Magpie.
@@ 2 Magpie: A Scalable Method to Synthesize Instruction Data
In what follows, we describe our method, Magpie, to synthesize instruction data for fine-tuning LLMs. An instance of instruction data consists of at least one or multiple instruction-response pairs. Each pair specifies the roles of instruction provider and follower, along with their instruction and response. As shown in Figure 1, Magpie consists of two steps: (1) instruction generation, and (2) response generation. The pipeline of Magpie can be fully automated without any human intervention. Given the data generated by Magpie, practitioners may customize and build their own personalized instruction dataset accordingly (see Section 3 and Appendix B for more details). We detail each step in the following.
Step 1: Instruction Generation.
@@
The goal of this step is to generate an instruction for each instance of instruction data. Given an open-weight aligned LLM (e.g., Llama-3-70B-Instruct), Magpie crafts an input query in the format of the predefined instruction template of the LLM. This query defines only the role of instruction provider (e.g., user), and does not provide any instruction. Note that the auto-regressive LLM has been fine-tuned using instruction data in the format of the predefined instruction template. Thus, the LLM autonomously generates an instruction when the query crafted by Magpie is given as an input. Magpie stops generating the instruction once the LLM produces an end-of-sequence token. Sending the crafted query to the LLM multiple times leads to a set of instructions. Compared with existing synthetic approaches [16, 31, 47, 53, 55, 58, 59], Magpie does not require specific prompt engineering techniques since the crafted query follows the format of the predefined instruction template. In addition, Magpie autonomously generates instructions without using any seed question, ensuring the diversity of generated instructions.
@@
Step 2: Response Generation. The goal of this step is to generate responses to the instructions obtained from Step 1. Magpie sends these instructions to the LLM to generate the corresponding responses. Combining the roles of instruction provider and follower, the instructions from Step 1, and the responses generated in Step 2 yields the instruction dataset. Detailed discussion on the generation configuration can be found in Appendix D.
Extensions of Magpie.
Magpie can be readily extended to generate multi-turn instruction datasets and preference datasets. In addition, practitioners can specify the task requested by the instructions. We defer the detailed discussion on these extensions to Appendix A.
@@ 3 Dataset Analysis
We apply Magpie to the Llama-3-8B-Instruct and Llama-3-70B-Instruct models to construct two instruction datasets: Magpie-Air and Magpie-Pro, respectively. Examples of instances in both datasets can be found in Appendix G. In this section, we present a comprehensive statistical analysis of the Magpie-Air and Magpie-Pro datasets. An overview of the lengths of instructions and responses of the data in Magpie-Air and Magpie-Pro is presented in Figure 3. In what follows, we first assess the breadth of Magpie-Pro by analyzing its coverage. We then discuss the attributes of Magpie-Pro, including topic coverage, difficulty, quality, and similarity of instructions, as well as quality of response. Finally, we provide the safety analysis and cost analysis. Using our dataset analysis, practitioners can customize and configure their own datasets for fine-tuning LLMs. In Appendix B, we showcase the process of customizing and filtering an instruction dataset based on our analysis. Specifically, we select 300K instances from Magpie-Pro and Magpie-Air-Filtered, yielding datasets Magpie-Pro-300K and Magpie-Air-300K-Filtered, respectively.
@@ 3.1 Dataset Coverage
We follow the approach in [64] and analyze the coverage of Magpie-Pro in the embedding space. Specifically, we use the all-mpnet-base-v2 embedding model [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 to calculate the input embeddings, and employ t-SNE [51] to project these embeddings into a two-dimensional space. We adopt three synthetic datasets as baselines, including Alpaca [47], Evol Instruct [58], and UltraChat [16], to demonstrate the coverage of Magpie-Pro.
Figure 3 presents the t-SNE plots of Magpie-Pro, Alpaca, Evol Instruct, and UltraChat. Each t-SNE plot is generated by randomly sampling 10,000 instructions from the associated dataset. We observe that the t-SNE plot of Magpie-Pro encompasses the area covered by the plots of Alpaca, Evol Instruct, and UltraChat. This suggests that Magpie-Pro provides a broader or more diverse range of topics, highlighting its extensive coverage across varied themes and subjects. We also follow the practice in [53] and present the most common verbs and their top direct noun objects in instructions in Appendix C, indicating the diverse topic coverage of Magpie dataset. Coverage analysis of Magpie-Air can also be found in Appendix C.
Figure 2: Lengths of instructions and responses in Magpie-Air/Pro.

### v2

@@ Abstract
High-quality instruction data is critical for aligning large language models (LLMs). Although some models, such as Llama-3-Instruct, have open weights, their alignment data remain private, which hinders the democratization of AI. High human labor costs and a limited, predefined scope for prompting prevent existing open-source data creation methods from scaling effectively, potentially limiting the diversity and quality of public alignment datasets. Is it possible to synthesize high-quality instruction data at scale by extracting it directly from an aligned LLM? We present a self-synthesis method for generating large-scale alignment data named Magpie. Our key observation is that aligned LLMs like Llama-3-Instruct can generate a user query when we input only the pre-query templates up to the position reserved for user messages, thanks to their auto-regressive nature. We use this method to prompt Llama-3-Instruct and generate 4 million instructions along with their corresponding responses. We further introduce extensions of Magpie for filtering, generating multi-turn, preference optimization, domain-specific and multilingual datasets. We perform a comprehensive analysis of the Magpie-generated data. To compare Magpie-generated data with other public instruction datasets (e.g., ShareGPT, WildChat, Evol-Instruct, UltraChat, OpenHermes, Tulu-V2-Mix, GenQA), we fine-tune Llama-3-8B-Base with each dataset and evaluate the performance of the fine-tuned models. Our results indicate that using Magpie for supervised fine-tuning (SFT) solely can surpass the performance of previous public datasets utilized for both SFT and preference optimization, such as direct preference optimization with UltraFeedback. We also show that in some tasks, models supervised fine-tuned with Magpie perform comparably to the official Llama-3-8B-Instruct, despite the latter being enhanced with 10 million data points through SFT and subsequent preference optimization. This advantage is evident on alignment benchmarks such as AlpacaEval, ArenaHard, and WildBench.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 (Achiam et al., 2023) and Llama-3 (Meta, 2024) have become integral to AI applications due to their exceptional performance on a wide array of tasks by following instructions. The success of LLMs is heavily reliant on the data used for instruction fine-tuning, which equips them to handle a diverse range of tasks, including those not encountered during training. The effectiveness of instruction tuning depends crucially on access to high-quality instruction datasets. However, the alignment datasets used for fine-tuning models like Llama-3-Instruct are typically private, even when the model weights are open, which impedes the democratization of AI and limits scientific research for understanding and enhancing LLM alignment.
@@
To address the challenges in constructing high-quality instruction datasets, researchers have developed two main approaches. The first type of method involves human effort to generate and curate instruction data (Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024; Zheng et al., 2024; 2023), which is both time-consuming and labor-intensive (Liu et al., 2024a). In contrast, the second type of methods uses LLMs to produce synthetic instructions (Ding et al., 2023; Yin et al., 2023; Li et al., 2024a; Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; 2024c; Xu et al., 2023a; b; Li et al., 2023a). Although these methods reduce human effort, its success heavily depends on prompt engineering and the careful selection of initial seed questions. The diversity of synthetic data tends to decrease as the dataset size grows. Despite ongoing efforts, the scalable creation of high-quality and diverse instruction datasets continues to be a challenging problem.
@@
Is it possible to synthesize high-quality instructions at scale by directly extracting data from advanced aligned LLMs? A typical input to an aligned LLM contains three key components: the pre-query template, the query, and the post-query template. For instance, an input to Llama-2-chat could be “[INST] Hi! [/INST]”, where [INST] is the pre-query template and [/INST] is the post-query template. These templates are predefined by the creators of the aligned LLMs to ensure the correct prompting of the models. We observe that when we only input the pre-query template to aligned LLMs such as Llama-3-Instruct, they self-synthesize a user query due to their auto-regressive nature. Our experiments indicate that these random user queries are of high quality and great diversity, suggesting that the abilities learned during the alignment process are effectively utilized.
Figure 1: This figure illustrates Magpie, the process of self-synthesizing alignment data from aligned LLMs (e.g., Llama-3-8B-Instruct) to create a high-quality instruction dataset. In Step 1, we input only the pre-query template into the aligned LLM and generate an instruction along with its response using auto-regressive generation. In Step 2, we use a combination of a post-query template and another pre-query template to wrap the instruction generated from Step 1, prompting the LLM to generate the response. This completes the construction of the instruction dataset. Magpie efficiently generates diverse and high-quality instruction data, which can be further extended to multi-turn (Magpie-MT), preference optimization (Magpie-DPO), domain-specific, and multilingual datasets.
@@
Based on these findings, we developed a self-synthesis method to construct high-quality instruction datasets at scale, named Magpie (as illustrated in Figure 1). Unlike existing methods, our approach does not rely on prompt engineering or seed questions. Instead, it directly constructs instruction data by prompting aligned LLMs with a pre-query template for sampling instructions. We also demonstrated the extensibility of Magpie in generating multi-turn, preference optimization, domain-specific, and multilingual datasets. We applied Magpie to the Llama-3-8B-Instruct and Llama-3-70B-Instruct models, creating two instruction datasets: Magpie-Air and Magpie-Pro, respectively.
@@
Our Magpie-Air and Magpie-Pro datasets were created using 206 and 614 GPU hours, respectively, without any human intervention or API access to production LLMs like GPT-4. The statistics and advantages of Magpie datasets compared to existing ones are summarized in Table 4 in Appendix A. We perform a comprehensive analysis of these two datasets in Section 3, allowing practitioners to filter and select data instances for fine-tuning models according to their particular needs.
@@
To compare Magpie data with other public instruction datasets (e.g., ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), Evol Instruct (Xu et al., 2023a), UltraChat (Ding et al., 2023), OpenHermes (Teknium, 2023a; b), GenQA (Chen et al., 2024), Tulu V2 Mix (Ivison et al., 2023)), we conducted supervised fine-tuning (SFT) of the Llama-3-8B-Base model with each dataset and assess the performance of the fine-tuned models on alignment benchmarks such as AlpacaEval 2 (Li et al., 2023b), Arena-Hard (Li et al., 2024b), and WildBench (Lin et al., 2024). Our results show that models supervised fine-tuned with Magpie achieve superior performance, even surpassing models that utilize both SFT and direct preference optimization (DPO) (Rafailov et al., 2023) with UltraFeedback (Cui et al., 2023). Notably, Magpie-aligned models outperform the official Llama-3-8B-Instruct model on AlpacaEval 2, despite the latter being fine-tuned with over 10 million data points for SFT and subsequent preference optimization. Not only does Magpie excel in SFT alone compared to prior public datasets, but also delivers the best results when combined with preference optimization methods such as DPO. By leveraging Magpie extensions to generate high-quality preference optimization datasets, Magpie-aligned Llama-3 models can even outperform GPT-4-Turbo(1106) on AlpacaEval 2. These findings show the exceptional quality of instruction data generated by Magpie, enabling it to outperform even the official, extensively optimized, and proprietary LLMs.
2 Magpie: A Scalable Method to Synthesize Alignment Data
Chat Templates of Aligned LLMs. For an aligned LLM (e.g., Llama-3-8B-Instruct), the input sequence can be represented as $x=T_{pre-query}\oplus q\oplus T_{post-query}$. Here, $q$ is the user query (e.g., "What material should I use to build a nest?"), while $T_{pre-query}$ and $T_{post-query}$ are pre-query and post-query templates. The pre-query template shows up before the user query, and the post-query template is defined as the conversation template between the user query and the LLM response. These templates are defined by the model provider to ensure the correct prompting. For example, for Llama-3-8B-Insturct model, $T_{pre-query}$ = <|start_header_id|>user<|end_header_id|>, and $T_{post-query}$ =<|eot_id|><|start_header_id|>assistant<|end_header_id|>.
2.1 Magpie Pipeline
@@
Overview of Magpie. In what follows, we describe our lightweight and effective method, Magpie, to synthesize alignment data from aligned LLMs. An instance of instruction data consists of at least one or multiple instruction-response pairs. Each pair specifies the roles of instruction provider (e.g., user) and follower (e.g., assistant), along with their instruction and response. As shown in Figure 1, Magpie consists of two steps: (1) instruction generation, and (2) response generation. The Magpie pipeline can be fully automated without any human intervention, and can be readily adapted for the generation of multi-turn, preference, and domain-specific datasets, as detailed in Section 2.2. We describe each step in the following.
@@
Step 1: Instruction Generation. The goal of this step is to generate an instruction for each instance of instruction data. Given an open-weight aligned LLM (e.g., Llama-3-70B-Instruct), Magpie crafts a pre-query template in the format of the predefined instruction template of the LLM. Note that the auto-regressive LLM has been fine-tuned using instruction data in the format of the pre-query template. Thus, the LLM autonomously generates an instruction when the pre-query template crafted by Magpie is given as an input. Magpie stops generating the instruction once the LLM produces an end-of-sequence token. Sending the crafted query to the LLM multiple times leads to a set of instructions. We note that compared with existing synthetic approaches (Ding et al., 2023; Li et al., 2024a; Taori et al., 2023; Wang et al., 2023; 2024c; Xu et al., 2023a; b), Magpie does not require specific prompt engineering techniques since the crafted query follows the format of the predefined instruction template. In addition, Magpie autonomously generates instructions without using any seed question, ensuring the diversity of generated instructions.
@@
Step 2: Response Generation. The goal of this step is to generate responses to the instructions obtained from Step 1. Magpie sends these instructions to the LLM to generate the corresponding responses. Combining the roles of instruction provider and follower, the instructions from Step 1, and the responses generated in Step 2 yields the instruction dataset.
Applicability of Magpie on Different LLMs. Magpie can be readily deployed to state-of-the-art open-weight language models including but not limited to Llama-3 (Meta, 2024), Llama-3.1 (Dubey et al., 2024), Qwen2 (Yang et al., 2024), Gemma-2 (Team et al., 2024), and Phi-3 (Abdin et al., 2024). Please refer to Appendix A for detailed support and corresponding datasets.
Remark. Magpie generates high-quality instructions even when the instruction loss is masked during alignment. We hypothesize that LLMs retain an implicit memorization of instruction distributions. We leave it as a potential future research problem.
2.2 Magpie Extensions
Dataset Filtering. Magpie allows practitioners to select instruction data from the raw dataset generated from the above two steps based on their needs. In Appendix C, we explores potential filter configurations with eight available metrics for users to customize their own Magpie datasets. We also provide 6 off-the-shelf filter configurations and discuss their performance in Appendix F.4.
Generating Multi-Turn Instruction Datasets. Magpie can be readily extended to generate multi-turn instruction datasets. To construct a multi-turn dataset (denoted as Magpie-MT), we initially follow Steps 1 and 2 to generate the first turn of instruction and response. For subsequent turns, we append the pre-query template to the end of the full prompt from the previous round of communication. We observe that the model may occasionally forget its role as the user, especially for the 8B model. To mitigate this, we employ a system prompt designed to control the behavior of the LLM and reinforce its awareness of the multi-round conversation context. The full prompt for building the instructions of Magpie-MT can be found in Figure 14 in Appendix G. We follow the procedure in Step 2 of Section 2.1 to generate responses to form the multi-turn instruction dataset.
Generating Preference Optimization Datasets. Leveraging the diverse and high-quality instructions produced by Magpie, we present a simple and effective method for generating preference optimization data, inspired by Meng et al. (2024) and Tran et al. (2023). We first select a small proportion of high-quality instructions from the raw dataset generated by the Magpie pipeline, ensuring diverse task categories. For each selected instruction, we sample responses from the aligned LLM $k$ times, using a temperature of $T<1$. We then employ a reward model (RM) to annotate scores for these responses. The response with the highest RM score is labeled as the chosen response, while the one with the lowest RM score is designated as the rejected response.
Generating Domain-Specific and Multilingual Datasets. In certain scenarios, users may wish to fine-tune LLMs using domain-specific or multilingual instruction data to enhance performance within specific domains or languages. To address this need, we introduce a lightweight method to control both the task category and the language of generated instructions. Our approach involves guiding LLMs through a tailored system prompt, specifying that the model is a chatbot designed for a particular domain and outlining the types of user queries it might encounter. Figure 2 shows how to control generating math and Chinese instructions using system prompts. More examples of system prompts designed to control the generation of code, translation, and multilingual instructions are illustrated in Figure 15 in Appendix G.
Furthermore, we note that domain-specific and multilingual instruction data can also be generated using models that are tailored to particular fields. Magpie demonstrates broad applicability beyond diverse chat models, extending to specialized code models (e.g., DeepSeek-Coder-V2 (Zhu et al., 2024)) and math models (e.g., Qwen2-Math-7B-Instruct (Yang et al., 2024)). By leveraging the unique strengths and specializations of different models, Magpie can create a rich and diverse corpus of instructions. Examples of Magpie-generated instructions from different domain-specific models and multilingual models are provided in Appendix H.
Figure 2: This figure illustrates how to control generating math and Chinese instructions using system prompts. The system prompt in Chinese translates to: "You are an AI assistant designed to provide accurate and concise answers to users’ questions in Chinese." The user’s response in Chinese translates to: "I want to learn about some basic concepts and historical development trends of AI."
@@ 3 Dataset Analysis
To demonstrate the effectiveness of Magpie compared with baseline methods for generating diverse high-quality alignment datasets, we apply Magpie to the Llama-3-8B-Instruct and Llama-3-70B-Instruct models (Meta, 2024) to construct two instruction datasets: Llama-3-Magpie-Air (hereafter referred to as Magpie-Air) and Llama-3-Magpie-Pro (hereafter referred to as Magpie-Pro), respectively. Examples of instances in both datasets can be found in Appendix H. In this section, we present a comprehensive analysis of the Magpie-Air and Magpie-Pro datasets, including topic coverage, difficulty, quality, similarity of instructions, and the quality of the responses. We also provide the safety analysis and cost analysis of Magpie.
@@ 3.1 Dataset Coverage
We follow Zhao et al. (2024) and analyze the coverage of Magpie-Pro in the embedding space. Specifically, we use the all-mpnet-base-v2 embedding model [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 to calculate the input embeddings, and employ t-SNE (Van der Maaten & Hinton, 2008) to project these embeddings into a two-dimensional space. We adopt three synthetic datasets as baselines, including Alpaca (Taori et al., 2023), Evol Instruct (Xu et al., 2023a), and UltraChat (Ding et al., 2023), to demonstrate the coverage of Magpie-Pro. The detailed analysis can be found in Appendix D.1. We observe that the t-SNE plot of Magpie-Pro encompasses the area covered by the plots of Alpaca, Evol Instruct, and UltraChat. This suggests that Magpie-Pro provides a broader or more diverse range of topics.
We also follow Wang et al. (2023) and present the most common verbs and their top direct noun objects in instructions in Appendix D, indicating the diverse topic coverage of Magpie dataset.